A research firm conducted a survey to determine the mean amount of money smokers spend on cigarettes during a day. A sample of 100 smokers revealed that the sample mean is N$ 5.24 and sample standard deviation is N$ 2.18 Assume that the sample was drawn from a normal population.

2.1 Find the point estimate of the population mean.

2.2 Determine the lower limit of the 95% conference interval for estimating the unknown population
1
Expert's answer
2020-11-02T20:32:09-0500

2.1. The sample mean $5.24 is a point estimate of the unknown population mean and is the best estimate, given the data.

2.2. The number of degrees of freedom are "df=100-1=99," and the significance level is "\alpha=0.05." The critical t-value for  "\alpha=0.05" and "df=99" degrees of freedom is "t_c=1.984217."

The 95% confidence for the population mean "\mu" is computed using the following expression


"CI=(\bar{x}-\dfrac{t_c\times s}{\sqrt{n}}, \bar{x}+\dfrac{t_c\times s}{\sqrt{n}})="

"=(5.24-\dfrac{1.984217\times 2.18}{\sqrt{100}}, 5.24+\dfrac{1.984217\times 2.18}{\sqrt{100}})="

"=(4.807, 5.673)"

Lower limit is 4.807.